# PEGASUS: Physically Enhanced Gaussian Splatting Simulation System for 6DOF Object Pose Dataset Generation

### v1

@@ Abstract
We introduce Physically Enhanced Gaussian Splatting Simulation System (PEGASUS) for 6DOF object pose dataset generation, a versatile dataset generator based on 3D Gaussian Splatting. Environment and object representations can be easily obtained using commodity cameras to reconstruct with Gaussian Splatting. PEGASUS allows the composition of new scenes by merging the respective underlying Gaussian Splatting point cloud of an environment with one or multiple objects. Leveraging a physics engine enables the simulation of natural object placement within a scene through interaction between meshes extracted for the objects and the environment. Consequently, an extensive amount of new scenes - static or dynamic - can be created by combining different environments and objects. By rendering scenes from various perspectives, diverse data points such as RGB images, depth maps, semantic masks, and 6DoF object poses can be extracted. Our study demonstrates that training on data generated by PEGASUS enables pose estimation networks to successfully transfer from synthetic data to real-world data. Moreover, we introduce the Ramen dataset, comprising 30 Japanese cup noodle items. This dataset includes spherical scans that captures images from both object hemisphere and the Gaussian Splatting reconstruction, making them compatible with PEGASUS.
Index Terms:
dataset generation, robotics, radiance fields, sim2real
Figure 1: Representative scenes generated by PEGASUS. By separately reconstructing objects and environment with Gaussian Splatting and connecting them to a physics engine a vast variety of scenes can be generated utilizing novel view synthesis. At each snapshot multiple data points such as RGB images, segmentation masks, depth maps, 2D/3D bounding boxes and object poses can be extracted.
22footnotetext: This work was conducted during an internship at the Industrial CPS Research Center, National Institute of Advanced Industrial Science and Technology
@@ I Introduction
Numerous Western nations are contending with the challenges posed by shifting demographics [1], particularly the reduction in their working-age population [2]. Japan is grappling with a similar issue, experiencing a decrease in its overall population, prompting substantial investments in robotics across multiple industries like healthcare, manufacturing, and agriculture to maintain workforce stability [4].
The focus of our research is on the development of robotic systems in the service sector to support personnel in retail. With Japan having one of the highest population densities, it boasts a considerable number of 24-hour convenience stores known as konbini [3]. Introducing robots into these compact retail spaces can prove to be a valuable asset for tasks like restocking products and efficiently managing inventory.
To apply deep learning approaches for object pose estimation most datasets are focused towards western style product. Examples of such datasets are YCB-V [11], HOPE [8], Fallen Things (FAT) [9] and LINEMOD [10]. For generating domain specific datasets the default approach would be to synthetically generate data with NDDS [6] or BlenderProc [5]. This approach allows for the straightforward insertion of modeled object assets into synthetic environments. However, a drawback is the presence of a domain gap when trained on synthetic generated dataset due to the lack of realism. We overcome this problem by using photorealistic rendering by using novel view synthesis. An additional drawback is the time-consuming necessity of creating detailed models for specific environments and objects. On the other hand assets for PEGASUS can be simply obtained by scanning real world objects and environments by reconstructing them with novel view synthesis methods such as 3D Gaussian Splatting (3DGS) [12]. By combining 3DGS assets a comprehensive dataset can be created to fine tune object pose estimation networks for desired operating environments.
@@
In this work, we introduce PEGASUS, a physically enhanced Gaussian Splatting simulation environment, designed to create innovative datasets for 6DoF object pose estimation by leveraging advanced novel view synthesis techniques. In this framework we separately create the environment and objects with Gaussian Splatting and simulate the interaction of both elements using a physics engine. In this matter we can render novel views for RGB images, semantic masks, depth maps and metadata such as the object pose and 2D/3D bounding boxes. By extracting the data in the BOP data format [30] it can be easily used to train pose estimation networks and other network types.
@@
Our experiments demonstrate that the pose estimation network, Deep Object Pose (DOPE) [42], when trained on PEGASUS-generated dataset, can operate a grasping task with an UR5 based on our dataset and successfully shows synthetic to real transfer.
The contribution of this paper are:
@@
  • PEGASUS: A dataset generation tool for photo-realistic 6DoF object pose estimation, utilizing 3D Gaussian Splatting. The tool’s code has been made open-source [^1] PEGASUS: https://github.com/meyerls/PEGASUS.
@@
  • PEGASET Dataset: A collection of scanned environments and 21 objects (YCB-V).
@@ II Related Work
We first briefly review the different types of representation used for radiance fields and discuss methods to generate synthetic datasets.
@@
Neural Radiance Fields (NeRF) [20] build a continuous implicit representation by optimizing a Multi-Layer Perceptron (MLP) through volumetric rendering. This process encodes information within the MLP weights, requiring network queries for every spatial point to extract color and density data. Consequently, modifying NeRF involves adjusting these weights. CLIP-NeRF [24] integrates CLIP [25] to manipulate the shape and appearance of NeRF by training a deformation network, which limits its suitability for rapid editing.
InstantNGP [21] employs a multi-resolution hash grid for spatial information storage. Scene modifications necessitate altering the grid structure, as demonstrated by NeRFShop [22]. However, NeRFShop’s volumetric manipulation, which involves interactive region selection and manual object deformation, is not conducive to automation.
ADOP [26] and VET [27] utilize point cloud-based radiance fields, enhancing points with neural features. Despite their high-quality rendering and fast inference, these methods still rely on Multi View Stereo (MVS) for internal geometry, which is time consuming.
3DGS [12] introduces a novel approach combining point cloud rendering with Gaussian splats. Their specialized differential Gaussian rasterization pipeline facilitates straightforward manipulation (transformation, insertion, deletion) of the underlying point cloud. This enables an easy and rapid combination of different reconstructions, a crucial aspect of PEGASUS.
@@
NeuralLabeling [29] offers to directly annotate Neural Radiance Field’s and precisely extract the underlying object structures. To create a large dataset it is still time consuming to generate a vast variety of scenes.
PEGASUS adopts a hybrid approach, combining the strengths of both synthetic and real-world datasets. By leveraging the modularity of synthetic data, it allows for the generation of new scenes through the combination of scanned objects and diverse environments, leading to a multitude of data points. Additionally, PEGASUS employs novel view synthesis techniques to render photorealistic scenes that are practically indistinguishable from real-world data.
@@ III Prerequisites
In this section we cover the prerequisites on 3D Gaussian Splatting (3DGS) and 6DoF manipulation on 3DGS point clouds.
@@ III-A Gaussian Splatting
The input for Gaussian Splatting comprises a set of images capturing a static scene or object, corresponding poses, camera intrinsics, and a sparse point cloud, which are typically generated using Structure from Motion (SfM) [19]. This sparse point cloud is then transformed into a more complex 3D Gaussian Splatting point cloud, denoted as $\mathbf{P}_{GS}=\{\mathbf{P}_{\mu},\mathbf{P}_{\Sigma},\mathbf{P}_{\alpha},\mathbf{P}_{f}\}$. Each point in this cloud, represented by $\mathbf{x}$ (also interpretable as the mean position $\mathbf{\mu}$ of the Gaussian), is associated with a covariance matrix $\Sigma$, an opacity value $\alpha$, and a set of spherical harmonic coefficients $\mathbf{f}$, which are used for directional appearance coloring.
The 3D Gaussians have to be projected onto the 2D image plane in order to optimize the parameters of the Gaussian Splatting point cloud. Therefore a differential tile-based gaussian rasterization pipeline proposed by [12] is utilized. Each Gaussian is characterized by
@@
where $\mathbf{\Sigma}$ is a full 3D covariance matrix defined in world space and centered at the point mean $\mathbf{\mu}$. By back projecting the 3D Gaussians onto the 2D image the covariance matrix in image space [16] is computed through:
@@
$\mathbf{W}$ is defined as the transformation matrix from world to camera space and $\mathbf{J}$ being the Jacobian of the affine approximation of the projective transformation.
@@
$\mathbf{c_{i}}$ and $\alpha_{i}$ are defined as the color and opacity of the $i$th Gaussian.
@@
The manipulation of Gaussian Splatting benefits from its underlying explicit representation. The appearance of the Gaussian point cloud $\mathbf{P}_{GS}$ is defined by the Gaussian’s $\mathbf{P}_{\Sigma}$, 3D mean value $\mathbf{P}_{\mu}$, opacity values $\mathbf{P}_{\alpha}$ and coefficients $\mathbf{P}_{f}$ of the spherical harmonics. For manipulating a Gaussian point cloud only $\mathbf{P}_{\mu}$, $\mathbf{P}_{\Sigma}$ and $\mathbf{P}_{f}$ are relevant as the scalar opacity values in $\mathbf{P}_{\alpha}$ do not change if a transformation is applied. For appling a transformation matrix $\mathbf{T}=[\mathbf{R}|\mathbf{t}]$ the translational and rotational part has to be considered separately.
The translational part $\mathbf{t}$ is a straightforward operation. This involves applying a translation vector $\mathbf{t}_{\Delta}=(x_{\Delta},y_{\Delta},z_{\Delta})^{\top}$ to the mean values $(x,y,z)^{\top}$ of the Gaussian’s. The remaining parts of the Gaussian point cloud (such as $\mathbf{P}_{\Sigma}$ and $\mathbf{P}_{f}$) are not affected.
Rotating a Gaussian Splatting point cloud is more complex. Here a rotation matrix $\mathbf{R}$ has to be applied on the points $\mathbf{P}_{\mu}$, the covariance matrices $\mathbf{P}_{\Sigma}$ and the coefficients $\mathbf{P}_{f}$ of the spherical harmonics. For $\mathbf{P}_{\mu}$ and $\mathbf{P}_{\Sigma}$ the rotation is directly applied on their corresponding point clouds. If we want to obtain the same view dependent effects for the scene as the unrotated one, a rotation has to be applied to the spherical harmonics coefficients.
Spherical harmonics play a pivotal role in the domain of computer graphics, especially for rendering view-dependent color representation relative to a viewing angles denoted by $\theta$ (azimuth) and $\phi$ (elevation). Broadly speaking, they are akin to a 2D Fourier series expansion on the two-dimensional surface but defined on the surface of a sphere. The spherical harmonic base functions, represented by $Y_{l}^{m}(\theta,\phi)$, act as the orthonormal basis functions.
The real-valued spherical harmonic function can be expressed as:
$c(\theta,\phi)=\sum_{l=0}^{\infty}\sum_{m=-l}^{l}f_{lm}Y_{lm}(\theta,\phi).$ (4)
Where $\mathbf{c}$ denotes the un-normalized color value for one color channel, and $f_{lm}$ is the expansion coefficient for degree $l$ and mode $m$.
To rotate the spherical harmonics, one does not directly apply the rotation to the basis function . Instead, the transformation is carried out on the coefficients [38].
For this purpose the Wigner D-matrix can be utilized [37, 38]. For a given rotation matrix $\mathbf{R}$, the rotated SH coefficients $f^{\prime}_{lm}$ can be expressed as:
$f^{\prime}_{lm}=\sum_{m^{\prime}=-l}^{l}D^{(l)}_{mm^{\prime}}(\mathbf{R})f_{lm^{\prime}}$ (5)
Where $D^{(l)}_{mm^{\prime}}(\mathbf{R})$ is the Wigner D-function for the desired rotation of degree $l$.
@@ IV Methodology
In this chapter, we introduce PEGASUS (Physical Enhanced Gaussian Splatting Simulation System), a system designed to generate novel, multi-modal datasets using Gaussian Splatting. This method allows for the extraction of multiple data points, as illustrated in Fig. 1.
@@ IV-A1 Base Environment
The base environment serves as the foundational construct for building the actual scene in PEGASUS. To create this, we recorded 10 different scenes using a DSLR camera, capturing 100 to 150 images for each scene. We then used Structure from Motion (SfM)[19] to recover the set poses $\mathcal{E}$ and a sparse point cloud $\mathbf{P}_{\text{sparse}}\in\mathbb{R}^{N_{\text{sparse}}\times 3}$.
For true-to-scale metric reconstructions of indoor scenes, we followed the CherryPicker [33] approach, placing an aruco marker in the scene and scaling both the poses and the reconstruction according to the recovered scale factor. For outdoor scenes, we manually computed the scale factor.
To correctly align and center the scene in space (assuming planar scenes), we employed RANSAC [36] to fit a plane into the sparse point cloud. This process oriented the scene so that the $\mathbf{z}$ vector pointed upwards, and the center of mass of the inlier points was positioned at the center of the model coordinate system.
Subsequently, we performed Gaussian Splatting of the scene using the default parameters suggested by 3DGS [12] over $30.000$ iterations. Towards the end of this process, we extracted the plain point cloud by obtaining the mean value of each Gaussian splat. To integrate with the physics engine, we needed to convert this point cloud into a mesh. For this purpose, we employed the alpha shape algorithm [35] as a mesh recovery technique, transforming the 3DGS-extracted points into a geometric mesh.
Figure 3: Gallery of data generated by PEGASUS. It shows scenes generated with 10 different base environments and an arbitrary combination of the 30 elements from the Ramen dataset and from the 21 objects from the YCB-V dataset.
@@ IV-A2 Gaussian Splatting Object
The photometric entity is crucial for rendering objects using Gaussian Splatting. We take a spherical, sparse, and metrically reconstructed SfM model of the object as input. The Gaussian Splatting process, using the same settings as in the base environment, creates this entity. This photometric entity can then be integrated into the simulation, allowing simultaneous rendering of both the base environment and the object.
@@
For the geometric entity, we start with the colored point cloud extracted from the Gaussian Splatting reconstruction. The point cloud is first cleaned to remove outliers, and then we use the alpha shape algorithm [35] for mesh reconstruction. To achieve a smoother surface, we apply Laplacian smoothing [34]. The geometric entity is stored as a low-polygon triangle mesh, optimizing the computational efficiency when simulating this mesh with the physics engine.

### v2

@@ Abstract
We introduce Physically Enhanced Gaussian Splatting Simulation System (PEGASUS) for 6DoF object pose dataset generation, a versatile dataset generator based on 3D Gaussian Splatting. Environment and object representations can be easily obtained using commodity cameras to reconstruct with Gaussian Splatting. PEGASUS allows the composition of new scenes by merging the respective underlying Gaussian Splatting point cloud of an environment with one or multiple objects. Leveraging a physics engine enables the simulation of natural object placement within a scene through interaction between meshes extracted for the objects and the environment. Consequently, an extensive amount of new scenes - static or dynamic - can be created by combining different environments and objects. By rendering scenes from various perspectives, diverse data points such as RGB images, depth maps, semantic masks, and 6DoF object poses can be extracted. Our study demonstrates that training on data generated by PEGASUS enables pose estimation networks to successfully transfer from synthetic data to real-world data. Moreover, we introduce the Ramen dataset, comprising 30 Japanese cup noodle items. This dataset includes spherical scans that capture images from both the object hemisphere and the Gaussian Splatting reconstruction, making them compatible with PEGASUS.
Figure 1: Representative scenes generated by PEGASUS. By separately reconstructing objects and environment with Gaussian Splatting and connecting them to a physics engine a vast variety of scenes can be generated utilizing novel view synthesis. At each snapshot, multiple data points such as RGB images, segmentation masks, depth maps, 2D/3D bounding boxes, and object poses can be extracted.
11footnotetext: Visual Computing Erlangen, Friedrich-Alexander-Universität Erlangen-Nürnberg-Fürth, Germany. E-Mail: lukas.meyer@fau.de22footnotetext: National Institute of Advanced Industrial Science and Technology, Tokyo, Japan22footnotetext: This work was conducted during an internship at the National Institute of Advanced Industrial Science and Technology.11footnotetext: This paper is one of the achievements of joint research with and is owned copyrighted material of ROBOT Industrial Basic Technology Collaborative Innovation Partnership. This research has been supported by the New Energy and Industrial Technology Development Organization (NEDO), under the project ID JPNP20016.
@@ I Introduction
Robotic manipulation in changing environments with various novel objects pose significant challenges. Tasks such as object detection, segmentation, and pose estimation require substantial training data to adapt to new scenarios. The creation of suitable datasets can be achieved either by annotating real-world data (e.g., YCB-V [1], HOPE [2], LINEMOD [3]) or by utilizing synthetic datasets (e.g., Fallen Things (FAT) [4]). NDDS [5] and BlenderProc [6] synthetically generate domain-specific datasets, facilitating the straightforward insertion of modeled object assets into synthetic environments. However, training on synthetically generated datasets may be compromised by a lack of realism and the time-consuming process of creating detailed models for specific environments and objects.
To address these limitations and reduce the reality gap, we make use of advanced novel view synthesis techniques. The rendering quality of novel view synthesis methods, such as Neural Radiance Fields (NeRF) [7] have steadily increased over the past years [8]. With 3D Gaussian Splatting (3DGS) [9], a real-time method using explicit representation, it becomes possible to create more realistic scenes. By leveraging this explicit representation, we can build a modular pipeline by combining multiple Gaussian splatting point clouds for dataset generation. This approach allows for the rapid and flexible generation of high-quality, realistic synthetic data, which can significantly enhance training and adaptation for robotic manipulation tasks in diverse environments.
In this work, we introduce PEGASUS, a physically enhanced Gaussian Splatting simulation environment, designed to create innovative datasets for 6DoF object pose estimation. By utilizing 3DGS, it allows us to close or at least narrow the gap between synthetic training data and real-world applications. Furthermore, it is very simple to generate custom assets: new models are obtained by scanning real-world objects and environments and reconstructing them in an automated pipeline. By combining 3DGS assets a comprehensive dataset can be created to fine-tune object pose estimation networks for desired operating environments.
@@
In PEGASUS, we separately create the environment and objects with Gaussian Splatting and simulate the interaction of both elements using a physics engine. In this matter, we can render novel views for RGB images, semantic masks, depth maps, and metadata such as the object pose and 2D/3D bounding boxes. By extracting the data in the Benchmark for 6D Object Pose Estimation (BOP) data format [10] it can be easily used to train pose estimation networks and other network types.
@@
Our experiments demonstrate that the pose estimation network, Deep Object Pose (DOPE) [11] when trained on PEGASUS-generated dataset, can operate a grasping task with Universal Robots UR5 based on our dataset and successfully shows synthetic to real transfer. In summary, we make the following contributions:
@@
  • PEGASUS: A dataset generation tool for photo-realistic 6DoF object pose estimation, utilizing 3D Gaussian Splatting. The tool’s code has been made open-source [^1] PEGASUS Code: https://github.com/meyerls/PEGASUS.
@@
  • PEGASET Dataset: A collection of scanned environments and 21 re-scanned objects from YCB-V.
@@ II Related Work
Neural Radiance Fields (NeRF) [7] build a continuous implicit representation by optimizing a Multi-Layer Perceptron (MLP) through volumetric rendering. This process encodes information within the MLP weights, requiring network queries for every spatial point to extract color and density data. Consequently, editing NeRFs involves retraining, which is computationally expensive.
InstantNGP [19] employs a multi-resolution hash grid for spatial information storage, resulting in much faster access and thus training and rendering time. However, scene modifications necessitate altering the grid structure, as demonstrated by NeRFShop [20]. NeRFShop enables volumetric manipulation, yet this involves interactive region selection and manual object deformation, and it is not conducive to automation. CLIP-NeRF [22] integrates CLIP [23] to manipulate the shape and appearance of NeRF by training a deformation network, which limits its suitability for rapid editing.
Alternatives are non-volumetric representations, based on surfaces or points. Such representations are not trained from scratch, but start with a point cloud reconstructed using standard structure-from-motion methods [18], SLAM or LIDAR. ADOP [24] and TRIPS [25] fall into this class, they utilize point cloud-based radiance fields, enhancing points with neural features. These approaches achieve high rendering quality and fast rendering time, but the optimization of the neural features and the rendering network is still time-consuming.
3D Gaussian Splatting (3DGS) [9] also starts with a reconstructed point cloud, but renders these as semi-transparent Gaussian splats. Such splats are a powerful rendering primitive, which makes it possible to optimize them directly, without requiring neural parameters and a final neural rendering network. 3DGS show high rendering quality at a very high rendering speed. The specialized differential Gaussian rasterization pipeline facilitates straightforward manipulation (transformation, insertion, deletion) of the underlying point cloud. This enables an easy and rapid combination of different reconstructions, a crucial aspect of PEGASUS, which is why we 3DGS as the basis for our approach.
@@
NeuralLabeling [27] offers to directly annotate Neural Radiance fields and precisely extract the underlying object structures. To create a large dataset it is still time-consuming to generate a vast variety of scenes.
PEGASUS adopts a hybrid approach, combining the strengths of both synthetic and real-world datasets. Leveraging the modularity of synthetic data, it allows for the generation of new scenes through the combination of scanned objects and diverse environments, leading to a multitude of data points. Additionally, PEGASUS employs novel view synthesis techniques to render photorealistic scenes that are practically indistinguishable from real-world data.
@@ III-A Gaussian Splatting
The input for Gaussian Splatting comprises a set of images capturing a static scene or object, corresponding poses, camera intrinsics, and a sparse point cloud, which are typically generated using Structure from Motion (SfM) [18]. This sparse point cloud is then transformed into a more complex 3D Gaussian Splatting point cloud, denoted as $\mathbf{P}_{GS}=\{\mathbf{P}_{\mathbf{\mu}},\mathbf{P}_{\Sigma},\mathbf{P}_{\alpha},\mathbf{P}_{\mathbf{f}}\}$. Each point in this cloud, represented by $\mathbf{x}$ (also interpretable as the mean position $\mathbf{\mu}$ of the Gaussian), is associated with a covariance matrix $\mathbf{\Sigma}$, an opacity value $\alpha$, and a set of spherical harmonic coefficients $\mathbf{f}$, which are used for directional appearance coloring.
The 3D Gaussians have to be projected onto the 2D image plane to optimize the parameters of the Gaussian Splatting point cloud. Therefore a differential tile-based Gaussian rasterization pipeline proposed by [9] is utilized. Each Gaussian is characterized by
@@
where $\mathbf{\Sigma}$ is a full 3D covariance matrix defined in world space and centered at the point means $\mathbf{\mu}$. By projecting the 3D Gaussians back onto the 2D image the covariance matrix in image space [15] is computed through:
@@
$\mathbf{W}$ is defined as the transformation matrix from world to camera space and $\mathbf{J}$ is the Jacobian of the affine approximation of the projective transformation.
@@
$\mathbf{c_{i}}$ and $\alpha_{i}$ are defined as color and opacity of the $i$th Gaussian.
@@
The manipulation of Gaussian Splatting benefits from its underlying explicit representation. The appearance of the Gaussian point cloud $\mathbf{P}_{GS}$ is defined by the Gaussian’s $\mathbf{P}_{\Sigma}$, 3D mean value $\mathbf{P}_{\mu}$, opacity values $\mathbf{P}_{\alpha}$ and coefficients $\mathbf{P}_{f}$ of the spherical harmonics. For manipulating a Gaussian point cloud only $\mathbf{P}_{\mu}$, $\mathbf{P}_{\Sigma}$ and $\mathbf{P}_{f}$ are relevant as the scalar opacity values in $\mathbf{P}_{\alpha}$ do not change if a transformation is applied. For applying a transformation matrix $\mathbf{T}=[\mathbf{R}|\mathbf{t}]$ the translational and rotational part has to be considered separately.
The translational part $\mathbf{t}$ is a straightforward operation. This involves applying a translation vector $\mathbf{t}_{\Delta}=(x_{\Delta},y_{\Delta},z_{\Delta})^{\top}$ to the mean values $\mathbf{x}=(x,y,z)^{\top}$ of the Gaussians. The remaining parts of the Gaussian point cloud (such as $\mathbf{P}_{\Sigma}$ and $\mathbf{P}_{f}$) are not affected.
Rotating a Gaussian Splatting point cloud is more complex. Here, a rotation matrix $\mathbf{R}$ has to be applied on the points $\mathbf{P}_{\mu}$, the covariance matrices $\mathbf{P}_{\Sigma}$ and the coefficients $\mathbf{P}_{f}$ of the spherical harmonics. For $\mathbf{P}_{\mu}$ and $\mathbf{P}_{\Sigma}$ the rotation is directly applied to their corresponding point clouds.
To also obtain the identical view-dependent effects as for the original scene one has to apply a rotation also to the spherical harmonics [35]. This is commonly done by not rotating the base functions of the spherical harmonics but rather their coefficients [35]. In [37] a method is proposed to decompose the rotation matrix $\mathbf{R}$ into its Euler angles and build a block diagonal sparse matrix [35] to rotate every band individually. Due to the rotational in-variance of spherical harmonics, this rotation is lossless. For a detailed explanation please refer to [37, 35].
@@ IV-A1 Base Environment
The base environment serves as the foundational construct for building the actual scene in PEGASUS. To create this, we recorded 9 different planar scenes using a DSLR camera, capturing between 100 and 150 images per scene. We then used SfM to recover the set of poses and a sparse point cloud $\mathbf{P}_{\text{sparse}}$.
The scenes are then automatically scaled to obtain a true-to-scale metric reconstruction by placing an ArUco marker in the scene  [30] and afterward aligning the planar area to the center so that the $\mathbf{z}$ vector points upwards.
Subsequently, we performed Gaussian Splatting on the scene using the default parameters suggested by 3DGS [9]. Towards the end of this process, we extracted the plain point cloud by obtaining the mean value of each Gaussian splat. To integrate with the physics engine, we converted this point cloud into a mesh. For this purpose, we employed the alpha shape algorithm [33] as a mesh recovery technique, transforming the 3DGS-extracted points into a geometric mesh.
Figure 3: Gallery of data generated by PEGASUS. It shows scenes generated with 9 different base environments and an arbitrary combination of the 30 elements from the Ramen dataset and from the 21 YCB objects [32] from the YCB-V dataset.
@@ IV-A2 Gaussian Splatting Object
The photometric entity is crucial for rendering objects using Gaussian Splatting. We start with a spherical, sparse, and metrically reconstructed Structure from Motion (SfM) model of the object as input. Using Gaussian Splatting with the same settings as in the base environment setup, we generate this photometric entity. It can then be integrated into the simulation, enabling simultaneous rendering of both the base environment and the object.
@@
For the geometric entity, we start with the colored point cloud extracted from the Gaussian Splatting reconstruction. The point cloud is initially cleaned by removing outliers, and then the alpha shape algorithm [33] is applied for mesh reconstruction. To achieve a smoother surface, we use Laplacian smoothing [31]. The geometric entity is stored as a low-polygon triangle mesh, optimizing computational efficiency when simulating this mesh with the physics engine.